OpenStudy (anonymous):

Solve 7/8 P-4=10 please show your work.????

OpenStudy (anonymous):

First add 4 to both sides. You should now have (7/8)P=14. Next, divide by 7/8 (take the reciprocal) and you have P=14(8/7) So, P=16